Hello,
What would be the formula for solving this problem?
mn/4 - y = f; solve for n
Thank you

Answer by stanbon(57226)

mn/4 - y = f; solve for n
-------------------
mn/4 = f+y
---
mn = 4(f+y)
---
n = [4(f+y)]/m
====================
Cheers,
Stan H.
==============